Show that a sequence in R2 is convergent if and only if it is bounded and all it's convergent subsequences have the same limit


1
Expert's answer
2021-04-14T10:12:37-0400

Solution:

Every convergent sequence is bounded.

Proof. Let "\left(x_{n}\right)" be a sequence which converges to "\ell" , and let "\epsilon=1" and note that "\epsilon>0" . By the definition of convergence of a sequence, there exists some "N \in \mathbb{N}" such that if "n \geq N" , then "\left|x_{n}-\ell\right|<\epsilon" . We compute

"\begin{aligned}

&\left|x_{n}-\ell\right|<\epsilon \\

& \Longrightarrow-\epsilon<x_{n}-\ell<\epsilon \\

\Longrightarrow &-\epsilon+\ell<x_{n}<\epsilon+\ell \\

\Longrightarrow & \ell-1<x_{n}<\ell+1

\end{aligned}"

We make the following two calculations.

"\begin{aligned}

\ell & \leq|\ell| \\

\Longrightarrow \ell+1 & \leq|\ell|+1

\end{aligned}"

and

"\begin{aligned}

-\ell & \leq|\ell| \\

\Longrightarrow-|\ell| & \leq \ell \\

\Longrightarrow-|\ell|-1 & \leq \ell-1

\end{aligned}"

Therefore, we can compute

"\begin{aligned}
-|\ell|-1 \leq \ell-1<x_{n}<\ell+1 \leq \mid \ell+1 \\
\Longrightarrow-|\ell|-1<x_{n}<\mid \ell+1 \\
\Longrightarrow\left|x_{n}\right|<|\ell|+1

\end{aligned}"

Let "M=\max \left\{\left|x_{1}\right|,\left|x_{2}\right|, \ldots,\left|x_{N-1}\right|,|\ell|+1\right\}" . Then given any "x_{n}" , if "n<N,\left|x_{n}\right| \leq M" , and if "n \geq N" , then

"\left|x_{n}\right|<|\ell|+1 \leq M".

Hence the sequence "\left(x_{n}\right)" is bounded by M.

Next: If a sequence converges then all subsequences converge and all convergent subsequences converge to the same limit. 

Proof: Let "\left\{a_{n}\right\}_{n \in \mathbb{N}}" be any convergent sequence. Denote the limit by "\ell" . Let "\left\{b_{n}\right\}_{n \in \mathbb{N}}" be any subsequence. Let "\varepsilon>0" be given. By the definition of convergence for "\left\{a_{n}\right\}_{n \in \mathbb{N}}" there exists "N \in \mathbb{N}" such that "\left|a_{n}-\ell\right|<\varepsilon" for all "n \geq N" . But this value N will also work for "\left\{b_{n}\right\}_{n \in \mathbb{N}}" . This is because if "n \geq N,\ then\ b_{n}=a_{m}" for some "m \geq n \geq N" and so "\left|b_{n}-\ell\right|=\left|a_{m}-\ell\right|<\varepsilon" . Thus "\left|b_{n}-\ell\right|<\varepsilon" for all "n \geq N"

as required.